# LLM-Based Program Generation for Triggering Numerical Inconsistencies Across Compilers

### v1

@@ Abstract.
Floating-point inconsistencies across compilers can undermine the reliability of numerical software. We present LLM4FP, the first framework that uses Large Language Models (LLMs) to generate floating-point programs specifically designed to trigger such inconsistencies. LLM4FP combines Grammar-Based Generation and Feedback-Based Mutation to produce diverse and valid programs. We evaluate LLM4FP across multiple compilers and optimization levels, measuring inconsistency rate, time cost, and program diversity. LLM4FP detects over twice as many inconsistencies compared to the state-of-the-art tool, Varity. Notably, most of the inconsistencies involve real-valued differences, rather than extreme values like NaN or infinities. LLM4FP also uncovers inconsistencies across a wider range of optimization levels, and finds the most mismatches between host and device compilers. These results show that LLM-guided program generation improves the detection of numerical inconsistencies.
copyright: none
@@ 1. Introduction
Floating-point arithmetic is widely used in numerical software, scientific computing, and simulations. However, reasoning about its behavior remains challenging. Small rounding errors, precision issues and different compiler optimizations can all lead to unexpected behaviors (DBL, 1991). For example, when the same program is compiled with different compilers and/or executed on different architectures such as CPUs and GPUs, developers may observe inconsistent results, since different compilers may generate different code for the same program (Markstein, 2008). Such inconsistencies are not trivial, and detecting and understanding them is essential for building reliable numerical software.
@@
To the best of our knowledge, Varity (Laguna, 2020) is the state-of-the-art randomized differential testing framework for generating floating-point programs that trigger numerical inconsistencies across compilers and/or optimization levels. However, Varity mainly relies on random program generation and does not make use of any domain knowledge or prior test results. Many generated programs fail to trigger any inconsistency, leading to low testing efficiency. This limits Varity’s ability to find more subtle compiler-specific numerical inconsistencies within a reasonable budget of test programs. In a parallel line of work, Large Language Models (LLMs) have recently shown abilities in generating programs that have valid syntax and rich semantics (Jiang et al., 2024). Several tools have begun using LLMs to generate test programs for compiler fuzzing (Xia et al., 2024; Yang et al., 2024; Munley et al., 2024; Gu, 2023). However, they mainly target general-purpose code generation, instead of focusing on floating-point programs. As a result, existing LLM-based testing approaches are not well-suited for exposing numerical inconsistencies.
@@
Figure 1. LLM4FP Overview
We evaluate LLM4FP on a set of compilers and optimization levels. We compare against Varity, the current state-of-the-art tool, as well as two LLM-based baselines: Direct-Prompt and Grammar-Guided. To measure effectiveness, we define the inconsistency rate as the ratio of output mismatches to the total number of pairwise comparisons performed across compilers and optimization levels. The results show that LLM4FP achieves an inconsistency rate of 26.56%, more than twice that of the state-of-the-art tool Varity. This includes 4,781 total inconsistencies from a testing budget of 1,000 generated programs. Compared to baseline LLM variants, LLM4FP improves both inconsistency detection and program diversity. Notably, over 92% of the inconsistencies involve normal and subnormal numerical values instead of extreme values such as NaN or infinities, which indicates that LLM4FP is particularly effective at exposing subtle numerical inconsistencies. We further analyze inconsistency patterns across compilers and optimization levels. LLM4FP finds more inconsistencies between host and device compilers, and at lower optimization levels such as O1 and O2, while Varity mostly detects inconsistencies only at the most aggressive level O3_fastmath.
@@
  • We conduct a comprehensive evaluation across multiple compilers and optimization levels. LLM4FP outperforms existing approaches in inconsistency detection rate while overcoming the program diversity loss from other LLM-based variants, and it is especially effective at finding subtle floating-point inconsistencies.
@@ 2.1. Overview
Figure 1 shows the overall workflow of our approach, LLM4FP, for discovering floating-point inconsistencies across compilers. The first stage is to select a strategy for generating candidate programs. We propose two complementary strategies: (1) Grammar-Based Generation, which uses a given grammar specification to guide the structure of the generated programs; and (2) Feedback-Based Mutation, which randomly chooses a previously generated program that triggered numerical inconsistencies across compilers as basis to generate more programs. Once the prompt is constructed, we invoke the LLM to generate a new floating-point C program. The next stage is the Compilation Driver, which prepares the program for execution on both CPUs and GPUs. Thus this stage handles C to CUDA code translation and compilation using a configurable set of compilers and optimization flags. Finally, Differential Testing executes the programs on both host and device to detect inconsistencies in the results. If the generated program triggers any numerical inconsistencies, then it is added into the “successful” program set for potential reuse in the Feedback-Based Mutation phase. This creates a feedback loop where successful programs help guide the LLM to generate other inconsistency-triggering programs.
@@ 2.2. Floating-Point Program Structure
Instead of directly prompting the LLM to generate numerical programs without any constraints, we specify (1) a high-level program structure to facilitate differential testing, and (2) a syntax grammar that specifies the internal structure of the numerical code to reduce the probability of generating invalid programs. Our approach uses the LLM’s flexibility while guiding it to generate floating-point programs with a consistent structure tailored for exposing numerical variations. LLM4FP adopts the high-level program structure and grammar first introduced in Varity (Laguna, 2020), which has been shown to be simple yet effective in guiding the generation of floating-point programs that expose numerical variations. The above also simplifies the comparison between LLM4FP and Varity in our evaluation in Section 3.
@@
The internal structure of the function compute is determined by a grammar, shown in Figure 2. The grammar guides the LLM to produce meaningful floating-point computations by capturing common patterns from high-performance computing (HPC) programs (Laguna, 2020). These patterns are likely to influence how compilers generate and optimize code. Specifically, the grammar specification allows to specify the floating-point precision level, either single or double precision, to use in the program. It also includes arithmetic expressions involving basic operators in ${+,,,/}\{$+,-,*,/\}, parentheses “()()” for grouping, and invoking functions from the standard C Math Library. The grammar also allows nested loops to reflect typical HPC control structures, as well as conditional branches. Additionally, the use of temporary floating-point variables is possible, which may store either scalar values or arrays.
@@
| <fp -type > <id > <assign -op >
<expression > ";"
@@ 2.3.2. Feedback-Based Mutation
The prompt first asks the LLM to change a given floating-point C program to create a new one that behaves differently. Next, the instruction includes: (1) the precision, double or single, to use for floating-point variables, (2) the high-level program structure, as described in Section 2.2, (3) guidelines to improve robustness and code quality, similar to those in the Grammar-Based strategy, (4) a list of mutation strategies to consider, including reordering or deeply nesting arithmetic expressions, changing numeric constants, introducing new control flow like nested loops or conditionals, using different Math Library functions, and inserting intermediate computations, (5) a concrete previously successful example which is randomly selected from the set of successful programs that triggered inconsistencies. Finally, we include the output formatting instruction at the end of the prompt as in the last strategy. In this way, the prompt encourages the LLM to preserve relevant patterns while generating novel variations.
2.4. Finding Compiler Inconsistencies
After generating a program, in the Compilation Driver stage, we prepare the generated program for execution on both CPUs and GPUs. Following the approach used in Varity, we convert the generated C program into CUDA by defining the compute function as a __global__ kernal, which is launched from the main function using a single block and a single thread. We then compile the C version using multiple host compilers, and the CUDA version using a GPU compiler, each under various optimization levels. Only binaries that compile successfully are passed to the next stage.
In the last stage, we perform differential testing (McKeeman, 1998) to check whether different compiler configurations produce inconsistent outputs for the same program and inputs. Specifically, for each program, we compare the results from the standard output for every pair of compilers under each optimization level. A numerical inconsistency is recorded when two outputs are not equal in their bitwise representations, i.e., the hexadecimal encoding of the floating-point result, such as when two 64-bit doubles yield different 16-character strings. This ensures that even small floating-point differences can be detected, which may arise from compiler implementation differences, precision trade-offs, or specific optimizations. While not all inconsistencies indicate compiler bugs, detecting them is still valuable for understanding numerical instability. If a program exhibits at least one such inconsistency across any compiler pair, it is added to the successful set. These programs are then reused by the Feedback-Based Mutation strategy to generate programs that are more likely to trigger inconsistencies in future iterations.
@@ 3.1.2. Optimization Levels
To align with Varity, we use six optimization levels across all compilers, listed in Table 1. These levels are designed to explore the impact of optimization aggressiveness and floating-point behavior triggering numerical inconsistencies. O0_nofma minimizes compiler-induced floating-point transformations and is the most compliant with the IEEE 754 standard (Markstein, 2008). It uses -O0 and explicitly disables Fused Multiply-Add (FMA), which often introduces more floating-point inconsistencies (Laguna, 2020). O0 uses -O0 with FMA enabled. O1, O2 and O3 use increasing levels of compiler optimizations. O3_fastmath uses -O3 with aggressive floating-point optimizations, i.e., -ffast-math in gcc/clang and --use_fast_math in nvcc, which may break IEEE 754 compliance. O3_fastmath provides the highest performance gain, but with the lowest numerical compliance.
@@ 3.1.3. Program Generation Budget and FP Precision
For each approach, we set a generation budget of 1,000 programs, each paired with a unique set of input values. Every program is compiled using supported combinations of compilers and optimization levels, and all resulting binaries are executed. Thus, for each approach there are a total of 18,000 program runs: 1,000 programs $\times$ 3 compiler pairs $\times$ 6 optimization levels. By default, the programs are configured to use double-precision floating-point arithmetic, i.e., FP64, but could be easily extended to other precisions such as single-precision, i.e., FP32.
@@ 3.1.4. Implementation Details
LLM4FP is implemented in Python, and uses the GPT-4.1-mini (OpenAI, 2025) LLM in the feedback loop to generate floating-point programs. GPT-4.1-mini is a variant of GPT-4 (OpenAI, 2023), and is the state-of-the-art in small model performance that provides a balance between intelligence, speed, and cost. We set temperature to 1.2 to increase randomness in token sampling and obtain more varied outputs. To reduce repetition, we set frequency_penalty to 0.5 to discourage the model from repeating frequently used tokens, and presence_penalty to 0.6 to encourage new patterns in the generated code. These values are informed by findings from prior work that systematically studies hyperparameter impacts on code generation quality and diversity (Arora et al., 2024). In addition, a fixed probability of 0.5 is used to randomly select either the Grammar-Based Generation or Feedback-Based Mutation strategy for each new program.
@@ 3.2.1. Baselines
Varity (Laguna, 2020)
A framework that triggers floating-point numerical differences in heterogeneous systems via random program generation. As discussed earlier Varity uses a subset of the C language grammar to generate well-formed C/CUDA floating-point programs, and performs differential testing to find numerical inconsistencies across compilers and/or optimization levels.
@@ 3.2.2. Metrics
To evaluate the diversity of all generated programs, we use two techniques: similarity scoring (Ren et al., 2020) and code clone detection (Ain et al., 2019). Measuring diversity is important to ensure that the LLM is producing structurally and semantically varied test programs, increasing the chance of exposing a wider range of floating-point inconsistencies.
@@
We first compute the CodeBLEU score (Ren et al., 2020), a widely-used metric for measuring code similarity. CodeBLEU absorbs the strength of the BLEU score (Papineni et al., 2002) in the n-gram match, and further incorporates code syntax via abstract syntax trees and code semantics via data-flow. A lower score indicates more syntactic and semantic variation, i.e., less similarity in generated programs. Specifically, we calculate the pairwise CodeBLEU score between all $N$ generated programs and report the average.
@@
In addition to CodeBLEU, we analyze code clones using the NiCad Clone Detector (Cordy and Roy, 2011) to detect redundant generation patterns that are structurally identical but superficially different, and thus unlike to reveal different numerical inconsistencies. NiCad supports the commonly used four types of code clones, Type-1 through Type-4, as originally proposed by (Roy and Cordy, 2007). In our analysis, we focus on Type-1, Type-2, and Type-2c code clones. Type-1 clones are identical code fragments except for differences in whitespace or comments. Type-2 clones allow syntactic similarity with variations in identifiers, literals, and types. Type-2c, a stricter subtype specific to NiCad, corresponds to structurally identical code with only consistent renaming of identifiers. We do not consider Type-3 (clones with inserted, deleted, or reordered statements), or Type-4 (semantic clones with similar behavior but different syntax), as floating-point programs are often computation-intensive and highly sensitive to small syntactic or semantic changes. Small modifications may lead to significantly different numerical behaviors, making such code fragments of interest for potentially triggering numerical inconsistencies.
@@
It is important to acknowledge that even programs not identified as clones may still expose the same numerical inconsistencies. To detect such cases, a more detailed analysis is necessary. Tools like pLiner (Guo et al., 2020) or Ciel (Miao et al., 2023) can be employed, since given a program that triggers a numerical inconsistency across different compilers or optimization levels, these tools pinpoint the specific lines or expressions responsible for the inconsistency. Using this information, programs can be grouped into equivalence classes based on the root causes of their numerical inconsistencies.
@@
Table 2. Comparing LLM4FP with baselines in terms of numerical inconsistencies and time cost ($h$:$m$:$s$). Incons. Rate = inconsistency rate; # Incons. = total number of inconsistencies.
Approach Incons. Rate # Incons. Time Cost
Varity 11.93% 2,147 00:30:42
Direct-Prompt 13.43% 2,417 03:23:40
Grammar-Guided 15.80% 2,844 03:16:44
LLM4FP 26.56% 4,781 03:22:00

### v2

# LLM4FP: LLM-Based Program Generation for Triggering Floating-Point Inconsistencies Across Compilers

@@ Abstract.
[…]
floating-point arithmetic, differential testing, random testing, Large Language Models
journalyear: 2025copyright: ccconference: Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis (SC Workshops ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3731599.3767362isbn: 979-8-4007-1871-7/2025/11ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Software reliabilityccs: Software and its engineering Compilersccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Parallel programming languages
@@ 1. Introduction
Floating-point arithmetic is widely used in numerical software, scientific computing, and scientific simulations. However, reasoning about its behavior remains challenging. Small rounding errors, precision issues and different compiler optimizations can all lead to unexpected behaviors (DBLP:journals/csur/Goldberg91). For example, when the same program is compiled with different compilers and/or executed on different architectures such as CPUs and GPUs, developers may observe inconsistent results, since different compilers may generate different code for the same program (DBLP:conf/dagstuhl/Markstein08). Such compiler-induced inconsistencies can lead to numerical bugs in widely used libraries (DBLP:conf/kbse/FrancoGR17), and the floating-point research community has highlighted the difficulty of ensuring numerical accuracy and reproducibility across heterogeneous computing systems (DBLP:conf/correctness/GopalakrishnanL21), with correctness in scientific computing recognized as a major challenge in High Performance Computing (HPC) (DBLP:journals/corr/abs-2312-15640).
@@
To the best of our knowledge, Varity (DBLP:conf/ipps/Laguna20) is the state-of-the-art randomized differential testing framework for generating floating-point programs that trigger numerical inconsistencies across compilers and/or optimization levels. However, Varity mainly relies on random program generation and does not make use of any domain knowledge or prior test results. Many generated programs fail to trigger any inconsistency, leading to low testing efficiency. This limits Varity’s ability to find more subtle compiler-specific numerical inconsistencies within a reasonable budget of test programs. In a parallel line of work, Large Language Models (LLMs) have recently shown abilities in generating programs that have valid syntax and rich semantics (DBLP:journals/corr/abs-2406-00515). Several tools have begun using LLMs to generate test programs for compiler fuzzing (DBLP:conf/icse/XiaPTP024; DBLP:journals/pacmpl/YangDLY0J024; DBLP:journals/fgcs/MunleyJC24; DBLP:conf/sigsoft/Gu23). However, they mainly target general-purpose code generation, instead of focusing on floating-point programs. As a result, existing LLM-based testing approaches are not well-suited for exposing numerical inconsistencies.
Figure 1. LLM4FP Overview
@@
[…]
@@
  • We conduct a comprehensive evaluation across multiple compilers and optimization levels. LLM4FP outperforms existing approaches in both inconsistency detection rate and program diversity.
@@ 2.1. Overview
Figure 1 shows the overall workflow of our approach, LLM4FP, for discovering floating-point inconsistencies across compilers. The first stage is to select a strategy for generating candidate programs. We propose two complementary strategies: (1) Grammar-Based Generation, which uses a given grammar specification to guide the structure of the generated programs; and (2) Feedback-Based Mutation, which randomly chooses a previously generated program that triggered numerical inconsistencies across compilers as basis to generate more programs. Once the prompt is constructed, we invoke the LLM to generate a new floating-point C program. The next stage is the Compilation Driver, which prepares the program for execution on both CPUs and GPUs. Thus this stage handles C to CUDA code translation and compilation using a configurable set of compilers and optimization flags. Finally, Differential Testing executes the programs on both host and device to detect inconsistencies in the results. If the generated program triggers any numerical inconsistencies, then it is added into the “successful” program set for potential reuse in the Feedback-Based Mutation phase. This creates a feedback loop in which successful programs help guide the LLM to generate other inconsistency-triggering programs.
@@ 2.2. Floating-Point Program Structure
Instead of directly prompting the LLM to generate numerical programs without any constraints, we specify (1) a high-level program structure to facilitate differential testing, and (2) a syntax grammar that specifies the internal structure of the numerical code to reduce the probability of generating invalid programs. Our approach uses the LLM’s flexibility while guiding it to generate floating-point programs with a consistent structure tailored for exposing numerical inconsistencies. LLM4FP adopts the high-level program structure and grammar first introduced in Varity (DBLP:conf/ipps/Laguna20), which has been shown to be simple yet effective in guiding the generation of floating-point programs that expose numerical inconsistencies. The above also simplifies the comparison between LLM4FP and Varity in our evaluation in Section 3.
@@
The internal structure of the function compute is determined by a grammar, shown in Figure 2. The grammar guides the LLM to produce meaningful floating-point computations by capturing common patterns from HPC programs (DBLP:conf/ipps/Laguna20). These patterns are likely to influence how compilers generate and optimize code. Specifically, the grammar specification allows to indicate the floating-point precision level, either single or double precision, to use in the program. It also includes arithmetic expressions involving basic operators in ${+,,,/}\{$+,-,*,/\}, parentheses “()()” for grouping, and invoking functions from the standard C Math Library. The grammar also allows nested loops to reflect typical HPC control structures, as well as conditional branches. Additionally, the use of temporary floating-point variables is possible, which may store either scalar values or arrays. Index-based array accesses are allowed. For example, LLM4FP can generate programs that contain code patterns such as for (i=0; i<N; i++) a[i] = ...;. Nested loops with array indexing are also supported.
@@
| <fp-type> <id> <assign-op> <expression> ";"
@@ 2.3.2. Feedback-Based Mutation
The prompt first asks the LLM to change a given floating-point C program to create a new one that behaves differently. Next, the instruction includes: (1) the precision, double or single, to use for floating-point variables, (2) the high-level program structure, as described in Section 2.2, (3) guidelines to improve robustness and code quality, similar to those in the Grammar-Based strategy, (4) a list of mutation strategies to consider, including reordering or deeply nesting arithmetic expressions, changing numeric constants, introducing new control flow like nested loops or conditionals, using different Math Library functions, and inserting intermediate computations, (5) a concrete previously successful example which is randomly selected from the set of successful programs that triggered inconsistencies. Finally, we include the output formatting instruction at the end of the prompt as in the last strategy. In this way, the prompt instructs the LLM to preserve key aspects of the original program, including its high-level structure, floating-point precision, and successful computational patterns, while generating novel variations.
2.4. Finding Floating-Point Inconsistencies
After generating a program, in the Compilation Driver stage, we prepare the generated program for execution on both CPUs and GPUs. Following the approach used in Varity, we convert the generated C program into CUDA by defining the compute function as a __global__ kernel, which is launched from the main function using a single block and a single thread. We then compile the C version using multiple host compilers, and the CUDA version using a GPU compiler, each under various optimization levels. Only binaries that compile successfully are passed to the next stage.
In the last stage, we perform differential testing (DBLP:journals/dtj/McKeeman98) to check whether different compiler configurations produce inconsistent outputs for the same program and inputs. Specifically, for each program, we compare the results from the standard output for every pair of compilers under each optimization level. A floating-point inconsistency is recorded when two outputs are not equal in their bitwise representations, i.e., the hexadecimal encoding of the floating-point result, such as when two 64-bit doubles yield different 16-character strings. This ensures that even small floating-point differences can be detected, which may arise from compiler implementation differences, precision trade-offs, or specific optimizations. While not all inconsistencies indicate compiler bugs, detecting them is still valuable for understanding numerical instability. If a program exhibits at least one such inconsistency across any compiler pair, it is added to the successful set. These programs are then reused by the Feedback-Based Mutation strategy to generate programs that are more likely to trigger inconsistencies in future iterations.
@@ 3.1.2. Optimization Levels
To align with Varity, we use six optimization levels across all compilers, listed in Table 1. These levels are designed to explore the impact of optimization aggressiveness and floating-point behavior triggering numerical inconsistencies. O0_nofma minimizes compiler-induced floating-point transformations and is the most compliant with the IEEE 754 standard (DBLP:conf/dagstuhl/Markstein08). It uses -O0 and explicitly disables Fused Multiply-Add (FMA), which often introduces more floating-point inconsistencies (DBLP:conf/ipps/Laguna20). O0 uses -O0 with FMA enabled. O1, O2 and O3 use increasing levels of compiler optimizations. O3_fastmath uses -O3 with aggressive floating-point optimizations, i.e., -ffast-math in gcc/clang and --use_fast_math in nvcc, which may break IEEE 754 compliance. O3_fastmath provides the highest performance gain, but with the lowest numerical compliance. We include this level to explore the effect of extreme optimizations on floating-point consistency, providing a contrast to more conservative levels.
@@ 3.1.3. Program Generation Budget and FP Precision
For each approach, we set a budget of 1,000 programs, each paired with a unique set of input values. Every program is compiled using supported combinations of compilers and optimization levels, and all resulting binaries are executed. Thus, for each approach there are a total of 18,000 program runs: 1,000 programs $\times$ 3 compiler pairs $\times$ 6 optimization levels. By default, the programs use double-precision floating-point arithmetic, i.e., FP64, but they could be easily extended to other precisions such as single-precision, i.e., FP32.
@@ 3.1.4. Implementation Details
LLM4FP is implemented in Python, and uses GPT-4 (DBLP:journals/corr/abs-2303-08774) in the feedback loop to generate floating-point programs since this model is the state-of-the-art for a wide range of NLP-based reasoning tasks (DBLP:journals/corr/abs-2303-12712). Specifically, we use the gpt-4.1-2025-04-14 version of the model. We set temperature to 1.2 to increase randomness in token sampling and obtain more varied outputs. To reduce repetition, we set frequency_penalty to 0.5 to discourage the model from repeating frequently used tokens, and presence_penalty to 0.6 to encourage new patterns in the generated code. These values are informed by findings from prior work that systematically studies hyperparameter impacts on code generation quality and diversity (DBLP:journals/corr/abs-2408-10577). In addition, each new program is generated using either the Grammar-Based Generation or the Feedback-Based Mutation strategy, with probabilities of 0.3 and 0.7, respectively.
@@ 3.2.1. Baselines
Varity (DBLP:conf/ipps/Laguna20)
A framework that triggers floating-point numerical differences in heterogeneous systems via random program generation. As discussed earlier, Varity uses a subset of the C language grammar to generate well-formed C/CUDA floating-point programs, and performs differential testing to find numerical inconsistencies across compilers and/or optimization levels.
@@ 3.2.2. Metrics
To evaluate the diversity of all generated programs, we use two techniques: similarity scoring (DBLP:journals/corr/abs-2009-10297) and code clone detection (DBLP:journals/access/AinBAAM19). Measuring diversity is important to ensure that the LLM is producing structurally and semantically varied test programs, increasing the chance of exposing a wider range of floating-point inconsistencies.
@@
We first compute the CodeBLEU score (DBLP:journals/corr/abs-2009-10297), a widely-used metric for measuring code similarity. CodeBLEU absorbs the strength of the BLEU score (DBLP:conf/acl/PapineniRWZ02) in the n-gram match, and further incorporates code syntax via abstract syntax trees and code semantics via data-flow. A lower score indicates more syntactic and semantic variation, i.e., less similarity between the generated programs. Specifically, we calculate the pairwise CodeBLEU score between all $N$ generated programs and report the average.
@@
In addition to CodeBLEU, we analyze the programs using the NiCad Clone Detector (DBLP:conf/iwpc/CordyR11a) to detect redundant generation patterns that are structurally identical but superficially different, and thus unlike to reveal different numerical inconsistencies. NiCad supports the commonly used four types of code clones, Type-1 through Type-4, as originally proposed by (roy2007survey). In our analysis, we focus on Type-1, Type-2, and Type-2c code clones. Type-1 clones are identical code fragments except for differences in whitespace or comments. Type-2 clones allow syntactic similarity with variations in identifiers, literals, and types. Type-2c, a stricter subtype specific to NiCad, corresponds to structurally identical code with only consistent renaming of identifiers. We do not consider Type-3 (clones with inserted, deleted, or reordered statements), or Type-4 (semantic clones with similar behavior but different syntax), as floating-point programs are often computation-intensive and highly sensitive to small syntactic or semantic changes. Small modifications may lead to significantly different numerical behaviors, making such code fragments of interest for potentially triggering numerical inconsistencies.
@@
It is important to acknowledge that even programs not identified as clones may still expose the same numerical inconsistencies. To detect such cases, a more detailed analysis is necessary. Tools like pLiner (DBLP:conf/sc/0007LR20) or Ciel (DBLP:conf/supercomputer/MiaoLR23) can be employed, since given a program that triggers a numerical inconsistency across different compilers or optimization levels, these tools pinpoint the specific lines or expressions responsible for the inconsistency. Using this information, programs can be grouped into equivalence classes based on the root causes of their numerical inconsistencies.
@@
Table 2. Comparing LLM4FP with baselines in terms of numerical inconsistencies, time cost ($h$:$m$:$s$), and program diversity. Incons. Rate = inconsistency rate; # Incons. = total number of inconsistencies. Lower CodeBLEU score is better.
Approach Incons. Rate # Incons. Time Cost CodeBLEU
Varity 11.93% 2,147 00:30:42 0.3581
Direct-Prompt 14.23% 2,562 04:16:37 0.3561
Grammar-Guided 16.47% 2,964 05:56:13 0.3442
LLM4FP 29.33% 5,280 05:37:42 0.2788